OpenStudy (anonymous):

Solve log3 27 – x = 5 Solve log3 27 – x = 5 @Mathematics

OpenStudy (anonymous):

what number do you raise 3 to, inorder to get 27?

OpenStudy (anonymous):

the base is 3, thats how the problem is

OpenStudy (anonymous):

right i understand, but 3^(what???)=27

OpenStudy (anonymous):

Ohhh! that would be 5

OpenStudy (anonymous):

no, it would not

OpenStudy (anonymous):

yes but that says what power to you raise (3) to to get (27) then you solve from there to get your x

OpenStudy (anonymous):

Ok it would be 3

OpenStudy (anonymous):

very nice

OpenStudy (anonymous):

now, simple solve the equation

OpenStudy (anonymous):

yes so 3-x=5

OpenStudy (anonymous):

Yeah but dont you have to use the power product property for this?

OpenStudy (anonymous):

no, just solve : 3-x=5

OpenStudy (lilg132):

3 - x = 5 x = 5 - 3 x = 2

OpenStudy (anonymous):

Ok, so can you put this step by step?

OpenStudy (anonymous):

why cant you do?

OpenStudy (lilg132):

did you understand how to work out the log3 27?

OpenStudy (anonymous):

Yeah but what happens to the 5?

OpenStudy (anonymous):

what? havent you been paying attention

OpenStudy (anonymous):

Yess I have, but Im still a bit confused.

OpenStudy (lilg132):

ok once you solve the log3 27 the answer to that is 3 so you replace log3 27 with 3 in the equation and you will get 3 - x = 5 now you have to rearrange the equation to get x so you bring 3 to the other side so your only left with x x = 5 - 3 x = 2

OpenStudy (anonymous):

ok so \[\log_{3}(27)-x=5 \] so what power do you rasise 3 to to get 27? 3 so now you have 3-x=5 solve for x -x=2 x=-2

OpenStudy (lilg132):

damn your right :p i totally missed out the minus x so the answer is -x = 2 x = -2

OpenStudy (lilg132):

do you still not understand mitosuki?

OpenStudy (anonymous):

I checked the answer but when I plug in the -2 it doesnt work

OpenStudy (anonymous):

It is -216.

OpenStudy (anonymous):

Yeah cause you put it this way: 3^5 = 27 -x

OpenStudy (anonymous):

did you plug \[\log_{3}(27) \] in as (log(27)/log(3)) because you calculator can only do log in base 10 so you have to use log base change formula

OpenStudy (anonymous):

so to plug it into a calculator it should look like this ((log(27)/log(3))-5 which equals -2

OpenStudy (anonymous):

Yeah but why -5?

OpenStudy (anonymous):

well look at your starting equation and i jsut solved for x first so i could plug it into my calculator

OpenStudy (anonymous):

Ok, thank you I understand it now :)

OpenStudy (anonymous):

yep glad i could help